At 700 K, Kc for the reaction H2(g) + I2(g) ⇌ 2HI(g) is 2.0 × 1019. What is Kc for HI(g) ⇌ 1/2H2(g) + 1/2I2(g)?
1
Expert's answer
2020-09-21T06:39:08-0400

H2(g) + I2(g) ⇌ 2HI(g) (1)

HI(g) ⇌ 1/2H2(g) + 1/2I2(g) (2)


Kc(1) = [HI]2 / [H2][I2]

Kc(2) = [H2]1/2[I2]1/2 / [HI]

Expressing Kc(2) by using Kc(1), we get

Kc(2) = (1/Kc(1))1/2 = (1 / 2.0*1019)1/2 = 2.2*10-10

Answer: 2.2*10-10